A data shows that people spent an average of $373 over the Christmas day in 2008. The following data show the dollar amount spent by a random sample of 16 people during the Christmas day in 2014: {427, 274, 370, 238, 222, 215, 360, 434, 387, 395, 323, 63, 415, 429, 377, 279}

(a) Test the hypothesis that that people spent less than an average of $373 over the Christmas day in 2014 using α = 0.05.

(b) Estimate the p-value for this test.
Answer #1

Solution:

x x2
427 182329
274 75076
370 136900
238 56644
222 49284
215 46225
360 129600
434 188356
387 149769
395 156025
323 104329
63 3969
415 172225
429 184041
377 142129
279 77841
--- ---
∑x=5208 ∑x2=1854742

The sample mean is

Mean = (x / n) )


=427+274+370+238+222+215+360+434+387+395+323+63+415+429+377+279/16

=5208/16

=325.5

The sample standard is S

  S =( x2 ) - (( x)2 / n ) n -1

=1854742-(5208)216/15

=1854742-1695204/15

=159538/15

=10635.8667

=103.1303

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :    = 373

Ha : < 373

Test statistic = t

= ( - ) / s / n

= (325 - 373) / 103 / 16

= −1.864

Test statistic = t = −1.864

P-value =0.0410

= 0.05  

P-value <

0.0410 < 0.05

Reject the null hypothesis .

There is sufficient evidence to suggest that